# Marito: Structuring and Building Open Multilingual Terminologies for South African NLP

### v1

@@ Abstract
The critical lack of structured terminological data for South Africa’s official languages hampers progress in multilingual NLP, despite the existence of numerous government and academic terminology lists. These valuable assets remain fragmented and locked in non-machine-readable formats, rendering them unusable for computational research and development. Marito [^1] https://www.dsfsi.co.za/za-marito/ addresses this challenge by systematically aggregating, cleaning, and standardising these scattered resources into open, interoperable datasets. We introduce the foundational Marito dataset, released under the equitable, Africa-centered NOODL framework. To demonstrate its immediate utility, we integrate the terminology into a Retrieval-Augmented Generation (RAG) pipeline. Experiments show substantial improvements in the accuracy and domain-specific consistency of English-to-Tshivenda machine translation for large language models. Marito provides a scalable foundation for developing robust and equitable NLP technologies, ensuring South Africa’s rich linguistic diversity is represented in the digital age.
Marito: Structuring and Building Open Multilingual Terminologies for South African NLP
@@ 1 Introduction
To bridge this critical gap, we introduce Marito [^2] Marito is a Xitsonga (TSO) word that means words. the South African curated Terminology, Lexicon, and Glossary Project. The mission of Marito is not to create new terminology from scratch, but to systematically aggregate, digitise, and standardise these scattered, publicly-funded terminological assets. By transforming them into interoperable, machine-readable formats, we unlock their potential for a new wave of linguistic and computational applications.
This paper presents the foundational work and initial release of the Marito project. Our primary contributions are threefold: First, we release the first version of the Marito dataset, a structured, multilingual terminology resource covering key domains for South African languages. Second, we release this dataset under the novel, Africa-centered Nwulite Obodo Open Data License (NOODL) to ensure equitable data governance and local benefit-sharing (Okorie and Omino, 2024). Third, we demonstrate the dataset’s immediate practical value by integrating it into a Retrieval-Augmented Generation (RAG) pipeline, which yields substantial improvements in machine translation accuracy and consistency for an English-to-Tshivenda language pair.
Ultimately, Marito provides both a practical resource and a scalable framework for fostering robust NLP and language technologies that reflect the rich linguistic diversity of South Africa.
@@ 2 Motivation
South African universities are also key actors in this landscape, developing linguistic resources in response to national policies that mandate the use of indigenous languages in higher education (of Arts and Culture, 2003; of Higher Education and Training, 2020). Language units at institutions like the University of KwaZulu-Natal, the University of Pretoria, and North-West University have produced valuable discipline-specific glossaries and corpora. However, these academic contributions often suffer from the same fate as government resources: they remain siloed within institutional repositories, lacking the standardisation and interoperability required for broad integration into NLP and AI ecosystems. The need for interventions like the Universities South Africa Community of Practice for African Languages (COPAL) highlights this persistent fragmentation, which a systematic project like Marito is designed to address.
The core motivation for Marito is to unlock the potential of these dormant linguistic assets. By systematically digitising (Taljard et al., 2022), structuring, and releasing these resources under equitable licenses (Okorie and Omino, 2024; Rajab et al., 2025) that adhere to FAIR principles (Wilkinson et al., 2016), we can directly enhance AI and NLP capabilities for South Africa’s indigenous languages. Properly structured terminologies can be ingested to fine-tune large language models (LLMs), improve machine translation, and power a new generation of inclusive technologies like AI-driven spell checkers and voice assistants. This, in turn, empowers linguists, educators, and innovators to build culturally relevant, domain-specific applications, from healthcare diagnostics in isiZulu to financial literacy tools in Setswana. The transition to open, machine-readable resources is therefore a critical step towards ensuring that South Africa’s languages not only survive but thrive in the digital era, fulfilling the multilingual promise of both government and higher education policies.
@@ 3 Methodology
The methodology for Marito is centered on the curation, standardisation, and dissemination of existing linguistic resources, rather than the creation of new terminology from scratch. Our approach systematically aggregates terminologies from disparate sources to enhance their accessibility and utility for linguistic research, education, and computational applications.
@@ 3.2 Domain Coverage
The scope of Marito is intentionally domain-agnostic, allowing for the inclusion of terminology lists from a wide array of fields. For instance, the DSAC lists encompass domains such as Information and Communication Technology (ICT), Mathematics, Finance, Health Sciences, and Parliamentary Procedure. The StatsSA collection provides comprehensive multilingual terminology for statistics. Similarly, the Open Educational Resource Term Bank (OERTB) project focused on developing African language terminologies for higher education across multiple disciplines (University of Pretoria, 2019; Taljard, 2015). This broad coverage ensures the dataset’s utility across diverse research and application contexts.
@@ 3.4 Data Curation and Structuring
Table 1: Overview of the datasets aggregated in the initial release of Marito (v0).
@@ 3.5 Data Release and Availability
The dataset is openly licensed (see Section 3.6) and accessible on multiple platforms, including GitHub [^4] https://github.com/dsfsi/za-marito/, Zenodo, and HuggingFace, to align with FAIR data principles (Wilkinson et al., 2016). We plan to provide both bulk download options and API access. A feedback and validation interface is also under development to enable community-driven refinement by linguists, translators, and other stakeholders. This approach supports a virtuous cycle of continuous improvement, ensuring the resource remains relevant and accurate over time.
@@ 3.6 Licensing under NOODL
Standard open licenses, while promoting reuse, often fail to address the power asymmetries and historical contexts inherent in community-generated data. To ensure equitable governance, Marito adopts the Nwulite Obodo Open Data License (NOODL), an African-centered framework designed to protect local agency and mandate fair benefit-sharing  Okorie and Omino (2024).
In contrast to generic licenses, NOODL differentiates access based on user context, mandates reinvestment from commercial use by entities outside developing regions, and includes provisions to reinforce community control. For Marito, this means:
@@ 4 Terminology Applications
The Marito datasets are not merely curated artifacts; they are critical assets for both computational evaluation and linguistic inquiry. Their primary applications fall into two key areas: providing a much-needed benchmark for multilingual NLP models and enabling deep analysis of language in a multilingual context.
@@ 4.1 A Benchmark for Multilingual NLP Evaluation
A significant bottleneck in developing NLP for African languages is the lack of standardized, domain-specific evaluation benchmarks (Adelani et al., 2023). Marito directly addresses this gap by providing gold-standard terminologies that can be used to rigorously assess model performance.
@@
Furthermore, the aligned multilingual terms enable the evaluation of multilingual word embeddings (Ruder et al., 2019; Upadhyay et al., 2016). By performing cluster analysis or measuring the cosine similarity of term-pairs (Almeida and Xexéo, 2019), researchers can probe how well semantically equivalent concepts are co-located within a shared vector space (Glavaš et al., 2019). This provides crucial insights into the quality of cross-lingual representations, particularly for low-resource languages. By serving as an evaluation resource, Marito supports the kind of participatory and community-centered benchmarking necessary for building truly useful technologies (Nekoto et al., 2020).
@@ 5 Improving Translations with Terminology lists and RAG
To demonstrate the practical value of the Marito terminologies, we conducted experiments to assess their impact on improving machine translation quality for a low-resource language pair. Despite advances in LLMs, their performance often degrades when translating domain-specific or rare terms, especially for languages with limited high-quality parallel corpora like South Africa’s (Zhong et al., 2024). This can lead to critical misinterpretations, such as confusing the term register in a mathematics context (ridzhisitara) versus an electoral one (redzhistara) in Tshivenda.
@@ 5.2 Experimental Conditions
    RAG with Semantic Terms: Key terms (nouns, verbs, adverbs) were extracted from the source text using spaCy’s en_core_web_sm model. These terms were used to retrieve relevant entries from the Marito vector store to augment the LLM’s prompt.
@@ 5.4.1 Quantitative Results
As shown in Table 2, the inclusion of a RAG pipeline with Marito terminologies leads to substantial improvements in translation quality across all metrics for both models and domains.
@@ 6 Future Directions and Call for Open Data
The expansion of Marito depends on the continued identification and integration of scattered terminological resources. As shown in Table 3, numerous valuable glossaries exist across South African institutions, but their accessibility varies dramatically, from openly licensed datasets like Unisa’s robotics glossary to web portals from Stellenbosch University that prohibit bulk download.
@@
This landscape illustrates the vital need for a structured, centralized effort like Marito. We acknowledge the institutional constraints that may lead to restrictive access, such as the need to track usage for funding reports. However, we advocate for a collective shift towards open, machine-readable formats under clear, permissive licenses. This not only aligns with Findable, Accessible, Interoperable, and Reusable (FAIR) principles but also empowers researchers, language practitioners, and developers by providing the legal and technical clarity needed to innovate. Ensuring that South Africa’s indigenous languages thrive in the digital age requires a concerted effort to make these foundational resources openly and sustainably available.
@@ 7 Conclusion
This paper introduced Marito, a project that directly confronts the critical scarcity of structured, machine-readable terminologies for South Africa’s official languages. By systematically aggregating, cleaning, and standardizing fragmented resources from government and academic sources, we have created a foundational, open-access dataset. Our adoption of the Africa-centered NOODL license further ensures that these resources are used in a manner that is equitable and benefits their communities of origin.
We have demonstrated the immediate, practical value of this structured terminology through RAG experiments, which yielded substantial improvements in English-to-Tshivenda machine translation accuracy and consistency. This result validates our core premise: that well-curated, accessible terminologies are not merely an academic exercise but are essential for enhancing the performance of language technologies for low-resource languages. Ultimately, Marito serves as both a valuable new resource and a call to action, providing a scalable foundation for developing more inclusive and capable NLP technologies that reflect the rich linguistic diversity of South Africa and the African continent.
@@ 8 Limitations
While Marito successfully structures existing terminological resources into more accessible formats, several limitations frame the scope of this work and offer avenues for future research.
@@
Secondly, the machine translation experiments, while promising, serve primarily as a proof of concept to demonstrate utility. Our evaluation was limited to English-to-Tshivenda translation in two specific domains. A more exhaustive evaluation is needed to assess the impact of Marito across all 11 official languages and on a wider array of NLP tasks, such as named entity recognition (NER) or cross-lingual information retrieval. Future work should benchmark performance on diverse tasks and languages to fully understand the resource’s capabilities and constraints, following community-driven evaluation standards (Nekoto et al., 2020).

### v2

# Mafoko: Structuring and Building Open Multilingual Terminologies for South African NLP

@@ Abstract
The critical lack of structured terminological data for South Africa’s official languages hampers progress in multilingual NLP, despite the existence of numerous government and academic terminology lists. These valuable assets remain fragmented and locked in non-machine-readable formats, rendering them unusable for computational research and development. Mafoko addresses this challenge by systematically aggregating, cleaning, and standardising these scattered resources into open, interoperable datasets. We introduce the foundational Mafoko dataset, released under the equitable, Africa-centered NOODL framework. To demonstrate its immediate utility, we integrate the terminology into a Retrieval-Augmented Generation (RAG) pipeline. Experiments show substantial improvements in the accuracy and domain-specific consistency of English-to-Tshivenda machine translation for large language models. Mafoko provides a scalable foundation for developing robust and equitable NLP technologies, ensuring South Africa’s rich linguistic diversity is represented in the digital age.
Mafoko: Structuring and Building Open Multilingual Terminologies for South African NLP
@@ 1 Introduction
To bridge this critical gap, we introduce Mafoko [^1] Mafoko is a Setswana (TSN) word that means words. the South African curated Terminology, Lexicon, and Glossary Project. The mission of Mafoko is not to create new terminology from scratch, but to systematically aggregate, digitise, and standardise these scattered, publicly-funded terminological assets. By transforming them into interoperable, machine-readable formats, we unlock their potential for a new wave of linguistic and computational applications.
This paper presents the foundational work and initial release of the Mafoko project. Our primary contributions are threefold: First, we release the first version of the Mafoko dataset, a structured, multilingual terminology resource covering key domains for South African languages. Second, we release this dataset under the novel, Africa-centered Nwulite Obodo Open Data License (NOODL) to ensure equitable data governance and local benefit-sharing (Okorie and Omino, 2024). Third, we demonstrate the dataset’s immediate practical value by integrating it into a Retrieval-Augmented Generation (RAG) pipeline, which yields substantial improvements in machine translation accuracy and consistency for an English-to-Tshivenda language pair.
Ultimately, Mafoko provides both a practical resource and a scalable framework for fostering robust NLP and language technologies that reflect the rich linguistic diversity of South Africa.
@@ 2 Motivation
South African universities are also key actors in this landscape, developing linguistic resources in response to national policies that mandate the use of indigenous languages in higher education (of Arts and Culture, 2003; of Higher Education and Training, 2020). Language units at institutions like the University of KwaZulu-Natal, the University of Pretoria, and North-West University have produced valuable discipline-specific glossaries and corpora. However, these academic contributions often suffer from the same fate as government resources: they remain siloed within institutional repositories, lacking the standardisation and interoperability required for broad integration into NLP and AI ecosystems. The need for interventions like the Universities South Africa Community of Practice for African Languages (COPAL) highlights this persistent fragmentation, which a systematic project like Mafoko is designed to address.
The core motivation for Mafoko is to unlock the potential of these dormant linguistic assets. By systematically digitising (Taljard et al., 2022), structuring, and releasing these resources under equitable licenses (Okorie and Omino, 2024; Rajab et al., 2025) that adhere to FAIR principles (Wilkinson et al., 2016), we can directly enhance AI and NLP capabilities for South Africa’s indigenous languages. Properly structured terminologies can be ingested to fine-tune large language models (LLMs), improve machine translation, and power a new generation of inclusive technologies like AI-driven spell checkers and voice assistants. This, in turn, empowers linguists, educators, and innovators to build culturally relevant, domain-specific applications, from healthcare diagnostics in isiZulu to financial literacy tools in Setswana. The transition to open, machine-readable resources is therefore a critical step towards ensuring that South Africa’s languages not only survive but thrive in the digital era, fulfilling the multilingual promise of both government and higher education policies.
@@ 3 Methodology
The methodology for Mafoko is centered on the curation, standardisation, and dissemination of existing linguistic resources, rather than the creation of new terminology from scratch. Our approach systematically aggregates terminologies from disparate sources to enhance their accessibility and utility for linguistic research, education, and computational applications.
@@ 3.2 Domain Coverage
The scope of Mafoko is intentionally domain-agnostic, allowing for the inclusion of terminology lists from a wide array of fields. For instance, the DSAC lists encompass domains such as Information and Communication Technology (ICT), Mathematics, Finance, Health Sciences, and Parliamentary Procedure. The StatsSA collection provides comprehensive multilingual terminology for statistics. Similarly, the Open Educational Resource Term Bank (OERTB) project focused on developing African language terminologies for higher education across multiple disciplines (University of Pretoria, 2019; Taljard, 2015). This broad coverage ensures the dataset’s utility across diverse research and application contexts.
@@ 3.4 Data Curation and Structuring
Table 1: Overview of the datasets aggregated in the initial release of Mafoko (v0).
@@ 3.5 Data Release and Availability
The dataset is openly licensed (see Section 3.6) and accessible on multiple platforms, including GitHub, Zenodo, and HuggingFace, to align with FAIR data principles (Wilkinson et al., 2016). We plan to provide both bulk download options and API access. A feedback and validation interface is also under development to enable community-driven refinement by linguists, translators, and other stakeholders. This approach supports a virtuous cycle of continuous improvement, ensuring the resource remains relevant and accurate over time.
@@ 3.6 Licensing under NOODL
Standard open licenses, while promoting reuse, often fail to address the power asymmetries and historical contexts inherent in community-generated data. To ensure equitable governance, Mafoko adopts the Nwulite Obodo Open Data License (NOODL), an African-centered framework designed to protect local agency and mandate fair benefit-sharing  Okorie and Omino (2024).
In contrast to generic licenses, NOODL differentiates access based on user context, mandates reinvestment from commercial use by entities outside developing regions, and includes provisions to reinforce community control. For Mafoko, this means:
@@ 4 Terminology Applications
The Mafoko datasets are not merely curated artifacts; they are critical assets for both computational evaluation and linguistic inquiry. Their primary applications fall into two key areas: providing a much-needed benchmark for multilingual NLP models and enabling deep analysis of language in a multilingual context.
@@ 4.1 A Benchmark for Multilingual NLP Evaluation
A significant bottleneck in developing NLP for African languages is the lack of standardized, domain-specific evaluation benchmarks (Adelani et al., 2023). Mafoko directly addresses this gap by providing gold-standard terminologies that can be used to rigorously assess model performance.
@@
Furthermore, the aligned multilingual terms enable the evaluation of multilingual word embeddings (Ruder et al., 2019; Upadhyay et al., 2016). By performing cluster analysis or measuring the cosine similarity of term-pairs (Almeida and Xexéo, 2019), researchers can probe how well semantically equivalent concepts are co-located within a shared vector space (Glavaš et al., 2019). This provides crucial insights into the quality of cross-lingual representations, particularly for low-resource languages. By serving as an evaluation resource, Mafoko supports the kind of participatory and community-centered benchmarking necessary for building truly useful technologies (Nekoto et al., 2020).
@@ 5 Improving Translations with Terminology lists and RAG
To demonstrate the practical value of the Mafoko terminologies, we conducted experiments to assess their impact on improving machine translation quality for a low-resource language pair. Despite advances in LLMs, their performance often degrades when translating domain-specific or rare terms, especially for languages with limited high-quality parallel corpora like South Africa’s (Zhong et al., 2024). This can lead to critical misinterpretations, such as confusing the term register in a mathematics context (ridzhisitara) versus an electoral one (redzhistara) in Tshivenda.
@@ 5.2 Experimental Conditions
    RAG with Semantic Terms: Key terms (nouns, verbs, adverbs) were extracted from the source text using spaCy’s en_core_web_sm model. These terms were used to retrieve relevant entries from the Mafoko vector store to augment the LLM’s prompt.
@@ 5.4.1 Quantitative Results
As shown in Table 2, the inclusion of a RAG pipeline with Mafoko terminologies leads to substantial improvements in translation quality across all metrics for both models and domains.
@@ 6 Future Directions and Call for Open Data
The expansion of Mafoko depends on the continued identification and integration of scattered terminological resources. As shown in Table 3, numerous valuable glossaries exist across South African institutions, but their accessibility varies dramatically, from openly licensed datasets like Unisa’s robotics glossary to web portals from Stellenbosch University that prohibit bulk download.
@@
This landscape illustrates the vital need for a structured, centralized effort like Mafoko. We acknowledge the institutional constraints that may lead to restrictive access, such as the need to track usage for funding reports. However, we advocate for a collective shift towards open, machine-readable formats under clear, permissive licenses. This not only aligns with Findable, Accessible, Interoperable, and Reusable (FAIR) principles but also empowers researchers, language practitioners, and developers by providing the legal and technical clarity needed to innovate. Ensuring that South Africa’s indigenous languages thrive in the digital age requires a concerted effort to make these foundational resources openly and sustainably available.
@@ 7 Conclusion
This paper introduced Mafoko, a project that directly confronts the critical scarcity of structured, machine-readable terminologies for South Africa’s official languages. By systematically aggregating, cleaning, and standardizing fragmented resources from government and academic sources, we have created a foundational, open-access dataset. Our adoption of the Africa-centered NOODL license further ensures that these resources are used in a manner that is equitable and benefits their communities of origin.
We have demonstrated the immediate, practical value of this structured terminology through RAG experiments, which yielded substantial improvements in English-to-Tshivenda machine translation accuracy and consistency. This result validates our core premise: that well-curated, accessible terminologies are not merely an academic exercise but are essential for enhancing the performance of language technologies for low-resource languages. Ultimately, Mafoko serves as both a valuable new resource and a call to action, providing a scalable foundation for developing more inclusive and capable NLP technologies that reflect the rich linguistic diversity of South Africa and the African continent.
@@ 8 Limitations
While Mafoko successfully structures existing terminological resources into more accessible formats, several limitations frame the scope of this work and offer avenues for future research.
@@
Secondly, the machine translation experiments, while promising, serve primarily as a proof of concept to demonstrate utility. Our evaluation was limited to English-to-Tshivenda translation in two specific domains. A more exhaustive evaluation is needed to assess the impact of Mafoko across all 11 official languages and on a wider array of NLP tasks, such as named entity recognition (NER) or cross-lingual information retrieval. Future work should benchmark performance on diverse tasks and languages to fully understand the resource’s capabilities and constraints, following community-driven evaluation standards (Nekoto et al., 2020).